OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x^2 - 6x - 7. Find f(g(2)). Can you explain me how to do that? Help, please?

OpenStudy (bibby):

start by finding g(2) plug the result into f(x)

OpenStudy (anonymous):

You mean g(2) = 2^2 - ^6(2) - 7?

OpenStudy (jhannybean):

Yes

OpenStudy (anonymous):

And then f(g(2)) = x(2^2 - ^6(2) - 7) + 8? Am I correct?

OpenStudy (bibby):

heck yeah

OpenStudy (anonymous):

Thank you :)

OpenStudy (jhannybean):

\[f(g(2)) = g(2) + 8\]

OpenStudy (bibby):

oh I didn't notice the x

OpenStudy (bibby):

there's also an extra ^ for some reason

OpenStudy (anonymous):

Eh, I am lost...?

OpenStudy (jhannybean):

Ok solve for g(2) fist. What do you get?

OpenStudy (anonymous):

It is 2^2 - ^(2) - 7?

OpenStudy (jhannybean):

I am not sure what `^(2)` means.

OpenStudy (anonymous):

it is 2 square?

OpenStudy (jhannybean):

\[g(2) = (2)^2 -6(2) -7 =~?\]

OpenStudy (anonymous):

Then it is 4 - 12 - 7 = -15?

OpenStudy (jhannybean):

yes.

OpenStudy (jhannybean):

Now.... \[f(g(2)) = -15 + 8=~?\]

OpenStudy (anonymous):

Um, but what about x + 8? But answer is -7

OpenStudy (jhannybean):

Exactly. What is -15 + 8?

OpenStudy (anonymous):

Oh, it is -7... Thank you :D

OpenStudy (jhannybean):

You're thinking too hard.

OpenStudy (anonymous):

Hahaha, sorry.